# LancBiO: Dynamic Lanczos-aided Bilevel Optimization via Krylov Subspace

### v1

@@ Appendix A Related Work in Bilevel Optimization
A variety of bilevel optimization algorithms are based on reformulation. These algorithms involve transforming the lower-level problem into a set of constraints, such as the optimal conditions of the lower-level problem (Dempe & Dutta, 2012; Li et al., 2023), or the optimal value condition (Outrata, 1990; Ye & Zhu, 1995, 2010; Dempe & Zemkoho, 2013; Lin et al., 2014; Xu & Ye, 2014). Furthermore, incorporating the constraints of the reformulated problem as the penalty function into the upper-level objective inspires a series of algorithms (Liu et al., 2022; Hu et al., 2023; Kwon et al., 2023a, b; Lu & Mei, 2023). Another category of methods in bilevel optimization is the iterative differentiation (ITD) based method (Maclaurin et al., 2015; Franceschi et al., 2017; Shaban et al., 2019; Grazzi et al., 2020; Liu et al., 2020; Ji et al., 2021), which takes advantage of the automatic differentiation technique. Central to this approach is the construction of a computational graph during each outer iteration, achieved by solving the lower-level problem. This setup facilitates the approximation of the hyper-gradient through backpropagation, and it is noted that parts of these methods share a unified structure, characterized by recursive equations (Ji et al., 2021; Li et al., 2022; Zhang et al., 2023). The approximate implicit differentiation (AID) treats the lower-level variable as a function of the upper-level variable. It calculates the hyper-gradient to implement alternating gradient descent between the two levels (Ghadimi & Wang, 2018; Ji et al., 2021; Chen et al., 2022; Dagréou et al., 2022; Li et al., 2022; Hong et al., 2023).
@@ Appendix B Krylov Subspace and Lanczos Process
Krylov subspace (Krylov, 1931) is fundamental in numerical linear algebra (Parlett, 1998; Saad, 2011; Golub & Van Loan, 2013) and nonlinear optimization (Yuan, 2014; Liu et al., 2021), specifically in the context of solving large linear systems and eigenvalue problems. We will briefly introduce the Krylov subspace and the Lanczos process, and recap some important properties; readers are referred to Saad (2011); Golub & Van Loan (2013) for more details.
@@
and the sequence of vectors $\left\{b,Ab,A^{2}b,\ldots,A^{N-1}b\right\}$ forms the basis for it. The Krylov subspace is widely acknowledged for its favorable properties in various aspects, including approximating eigenvalues (Kuczyński & Woźniakowski, 1992), solving the regularized nonconvex quadratic problems (Gould et al., 1999; Zhang et al., 2017; Carmon & Duchi, 2018), and reducing computation cost (Brown & Saad, 1990; Bellavia & Morini, 2001; Liu et al., 2013; Jiang et al., 2024).
The Lanczos process (Lanczos, 1950) is an algorithm that exploits the structure of the Krylov subspace when $A$ is symmetric. Specifically, in the $j$-th step of the Lanczos process, we can efficiently maintain an orthogonal basis $Q_{j}$ of $\mathcal{K}_{j}(A,b)$, so that $T_{j}=Q_{j}^{\top}AQ_{j}$ is tridiagonal, which means a tridiagonal matrix $T_{j}$ approximates $A$ in the Krylov subspace. Consequently, it allows to solve the minimal residual problem or the eigenvalue problem efficiently within the Krylov subspace. There are several equivalent variants of the Lanczos process (Paige, 1971, 1976; Meurant & Strakoš, 2006), and we follow the update rule as shown in Algorithm 3.
@@
We now present several key properties of the Krylov subspace and the Lanczos Process from Saad (2011).
@@
$v=-\frac{1}{c_{0}}\left(A^{m-1}v+c_{m-1}A^{m-2}v+\cdots+c_{2}Av+c_{1}v\right).$
@@
Denote the $n\times j$ matrix with column vectors $q_{1},\ldots,q_{j}$ by $Q_{j}$ and the $j\times m$ tridiagonal matrix by $T_{j}$, all of which are generated by Algorithm 3. Then the following 3-term recurrence holds.
@@
Based on Proposition B.4, the Lanczos process is illustrated in Figure 7
[…]
Figure 7: Classic 3-term recurrence
@@ Appendix C Dynamic Lanczos Subroutine
This section lists the DLanczos subroutine (Algorithm 4) invoked in the LancBiO framework (Algorithm 2). One of the main differences between Algorithm 3 and Algorithm 4 is that Algorithm 3 represents the entire $m$-step Lanczos process, while Algorithm 4 serves as a one-step subroutine. Specifically, LancBiO invokes DLanczos once in each outer iteration (line 11 in Algorithm 2), expanding both $T$ and $Q$ by one dimension. Consequently, the inputs of Algorithm 4 are not indexed to avoid confusion, with their corresponding variables $(T_{k-1},Q_{k-1},A_{k},\beta_{k})$ in Algorithm 2 evolving across outer iterations indexed by $k$. Another difference lies in the dynamic property of the DLanczos subroutine, i.e., the matrix $A_{k}$ passed during each invocation in Algorithm 2 varies, while the classic Lanczos process (Algorithm 3) employs a static $A$.
@@ Appendix D Extending LancBiO to Non-convex Lower-level Problem
The Lanczos process is known for its efficiency of constructing Krylov subspaces and is capable of solving indefinite linear systems (Greenbaum et al., 1999). In this section, we will briefly demonstrate that the dynamic Lanczos-aided Bilevel Optimization framework, LancBiO, can also handle lower-level problems with the indefinite Hessian.
@@
with initial ponit $x_{0}$, initial residual $r_{0}=b-Ax_{0}$ and initial error $e_{0}=A^{-1}b-x_{0}$. If the matrix $A$ is positive-definite, the classic Lanczos algorithm is equivalent to the Conjugate Gradient (CG) algorithm (Hestenes et al., 1952), both of which minimize the $A$-norm of the error in an affine space (Greenbaum, 1997; Meurant & Strakoš, 2006), i.e., at the $m$-th step,
@@
If the matrix $A$ is not positive-definite, MINRES (Paige & Saunders, 1975) is the algorithm recognized to minimize the 2-norm of the residual in an affine space (Greenbaum, 1997; Meurant & Strakoš, 2006), i.e., at the $m$-th step,
@@
Additionally, based on $Q_{m}$ as the basis of the Krylov subspace $\mathcal{K}_{m}(A,b)$, $T_{m}$ as the projection of $A$ onto $\mathcal{K}_{m}(A,b)$, and the 3-term recurrence
@@ Appendix E Proof of Smoothness of $y^{*}$ and $\varphi$
To ensure completeness, in this subsection, we provide detailed proofs for the preliminary lemmas that characterize the smoothness of the lower level solution $y^{*}$ and the hyper-objective $\varphi$.
@@
where $L_{\varphi}=\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left(L_{fx}+\frac{L_{gx}L_{fy}+L_{gxy}C_{fy}}{\mu_{g}}+\frac{L_{gx}C_{fy}L_{gyy}}{\mu_{g}^{2}}\right)$.
@@
$\left\|\left(A_{1}^{*}\right)^{-1}b_{1}-\left(A_{2}^{*}\right)^{-1}b_{2}\right\|\leq\frac{L_{fy}}{\mu_{g}}\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|+\frac{C_{fy}L_{gyy}}{\mu_{g}^{2}}\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|.$
In a similar way, we can derive
@@
$\left\|\nabla\varphi\left(x_{1}\right)-\nabla\varphi\left(x_{2}\right)\right\|\leq$ $~{}L_{fx}\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|$
$+L_{gx}\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left(\frac{L_{fy}}{\mu_{g}}+\frac{C_{fy}L_{gyy}}{\mu_{g}^{2}}\right)\left\|x_{1}-x_{2}\right\|$
$+L_{gxy}\frac{C_{fy}}{\mu_{g}}\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|$
$\leq$ $~{}L_{\varphi}\left\|x_{1}-x_{2}\right\|,$
where $L_{\varphi}=\left(1+\frac{L_{gy}}{\mu_{g}}\right)\left(L_{fx}+\frac{L_{gx}L_{fy}+L_{gxy}C_{fy}}{\mu_{g}}+\frac{L_{gx}C_{fy}L_{gyy}}{\mu_{g}^{2}}\right)$. ∎
@@
We rewrite dynamic update rule from Section 2.2
@@
Note that $Q_{1}=\mathrm{span}\{q_{1}\}$ with $q_{1}=\frac{\bar{b}}{\left\|{\bar{b}}\right\|}$ satisfies (19). We will give a proof by induction. Suppose for $i=1,2,\ldots,j$,
@@
Although we can estimate the difference between the basis of the above two subspaces, it is noted that the Krylov subspaces can be very sensitive to small perturbation (Meurant & Strakoš, 2006). Therefore, a more detailed observation of the dynamic Lanczos process is necessary, together with the development of relevant lemmas adapted from existing work (Paige, 1976, 1980; Greenbaum, 1997) to support our analysis.
@@
Then, we will consider the orthogonality of matrix $Q_{j}$, which is reflected by $R_{j}$ in (22). Multiply Equation 21 on the left by $Q^{\top}_{j}$,
@@
Combine its symmetry with the decomposition (22), then
@@
Denote $M_{j}=T_{j}R_{j}-R_{j}T_{j}$ which is upper triangular. Since the consecutive $q_{i}$ is orthogonal as revealed by (24), we can conclude that the diagnoal elements of $M_{j}$ are 00. Furthermore, by extracting the upper triangular part of the right hand side of (26), we can get
@@
From the bound of $\left\|{\delta q_{j}}\right\|$ it follows that for $1\leq s<t\leq j$, $\left|{\zeta_{st}}\right|\leq 2L_{gyy}\varepsilon_{j}$. ∎
Lemma F.2 illustrates the influence of the dynamics in Algorithm 2 imposed on the standard 3-term Lanczos recurrence, and as (22) reveals, $R$ characterizes the loss of orthogonality of the basis $Q$. However, the following lemmas demonstrate that the range of eigenvalues of the approximate projection matrix $T$ is controllable.
To begin with, we establish the Ritz pairs of $T_{j}$ as $\left(\mu_{i}^{\left(j\right)},y_{i}^{\left(j\right)}\right)$ for $i=1,2,\dots,j$, such that
@@
The subsequent lemma presents a description for the difference of eigenvalues between $T_{j}$ and the tridiagonal matrix in the preceding steps.
@@
Multiply the eigenvector $y_{r}^{\left(i\right)}$ of $T_{i}$ by $T_{j}$, where $i<j$,
$T_{j}\left[\begin{array}[]{c}y_{r}^{\left(i\right)}\\ \mathbf{0}\\ \end{array}\right]=\left[\begin{array}[]{c}\mu_{r}^{\left(i\right)}y_{r}^{\left(i\right)}\\ \beta_{r+1}\varsigma_{ir}^{\left(i\right)}e_{1}\\ \end{array}\right].$ (28)
@@
and multiply the eigenvectors $\left(y_{s}^{\left(j\right)}\right)^{\top}$ and $y_{t}^{\left(j\right)}$ on the left and right of (23), respectively,
@@
Note that $\left|\epsilon_{st}^{\left(j\right)}\right|\leq 2jL_{gyy}\varepsilon_{j}$ because of F.2. Specifically, by taking $s=t$ in (30),
@@
and we can rewrite (32) in the matrix form
@@
and it follows from the Cauchy–Schwarz inequality and (38),
$\min_{1\leq d\leq l<j}\left|{\mu_{i}^{\left(j\right)}-\mu_{d}^{\left(l\right)}}\right|\leq\frac{\sqrt{j}\sqrt{\sum_{r=1}^{j-1}{\sum_{t=1}^{r}{\left(\epsilon_{tt}^{\left(r\right)}\right)^{2}}}}}{\left|\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}\,\,\leq\frac{\sqrt{j}\sqrt{\sum_{r=1}^{j-1}{r^{2}\left(2L_{gyy}\varepsilon_{j}\right)^{2}}}}{\left|\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}\leq\frac{2j^{2}L_{gyy}\varepsilon_{j}}{\sqrt{3}\left|\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}.$
@@
Suppose Assumptions 3.1 to 3.3 hold. $Q_{j}$ and $T_{j}$ are the basis matrix and the approximate tridiagonal matrix in the $j$-th step and $R_{j}$ is the strictly upper triangular matrix defined in (22) characterizing the orthogonality of $Q_{j}$. Suppose $\mu_{i}^{\left(j\right)}$ is the $i$-th eigenvalue of $T_{j}$, then for $i=1,\ldots,j$,
@@
By conducting the left multiplication with $\left(y_{i}^{\left(j\right)}Q_{j}\right)^{\top}$ and the right multiplication with $y_{i}^{\left(j\right)}$ on both sides of equation (21), we obtain the following.
@@
Divide by $\left(z_{i}^{\left(j\right)}\right)^{\top}z_{i}^{\left(j\right)}$,
$\left|\frac{\left(z_{i}^{\left(j\right)}\right)^{\top}A_{j}^{*}z_{i}^{\left(j\right)}-\mu_{i}^{\left(j\right)}\left(z_{i}^{\left(j\right)}\right)^{\top}z_{i}^{\left(j\right)}}{\left(z_{i}^{\left(j\right)}\right)^{\top}z_{i}^{\left(j\right)}}\right|\leq\frac{\left|\epsilon_{ii}^{\left(j\right)}\right|+\left|\left(z_{i}^{\left(j\right)}\right)^{\top}\delta Q_{j}y_{i}^{\left(j\right)}\right|}{\left|1+2\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}\leq\frac{\left|\epsilon_{ii}^{\left(j\right)}\right|+L_{gyy}\varepsilon_{j}\sqrt{j\left(\left|1+2\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|\right)}}{\left|1+2\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|},$ (40)
since
@@
$\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}<\frac{3}{8}-\frac{\varepsilon_{j}}{2}$ (42)

### v2

@@ Appendix A Related Work in Bilevel Optimization
A variety of bilevel optimization algorithms are based on reformulation [Liu et al., 2021b, Yao et al., 2024]. These algorithms involve transforming the lower-level problem into a set of constraints, such as the optimal conditions of the lower-level problem [Dempe and Dutta, 2012, Li et al., 2023], or the optimal value condition [Outrata, 1990, Ye and Zhu, 1995, 2010, Dempe and Zemkoho, 2013, Lin et al., 2014, Xu and Ye, 2014]. Furthermore, incorporating the constraints of the reformulated problem as the penalty function into the upper-level objective inspires a series of algorithms [Liu et al., 2022, Hu et al., 2023, Kwon et al., 2023, Lu and Mei, 2024, Kwon et al., 2024]. Another category of methods in bilevel optimization is the iterative differentiation (ITD) based method [Maclaurin et al., 2015, Franceschi et al., 2017, Shaban et al., 2019, Grazzi et al., 2020, Liu et al., 2020, Ji et al., 2021], which takes advantage of the automatic differentiation technique. Central to this approach is the construction of a computational graph during each outer iteration, achieved by solving the lower-level problem. This setup facilitates the approximation of the hyper-gradient through backpropagation, and it is noted that parts of these methods share a unified structure, characterized by recursive equations [Ji et al., 2021, Li et al., 2022, Zhang et al., 2024]. The approximate implicit differentiation (AID) treats the lower-level variable as a function of the upper-level variable. It calculates the hyper-gradient to implement alternating gradient descent between the two levels [Ghadimi and Wang, 2018, Ji et al., 2021, Chen et al., 2022, Dagréou et al., 2022, Li et al., 2022, Hong et al., 2023]. Moreover, extending the spirits of AID to distributed settings has garnered increasing interest in recent years[Kong et al., 2024, He et al., 2024, Zhu et al., 2024].
@@ Appendix B Krylov Subspace and Lanczos Process
Krylov subspace [Krylov, 1931] is fundamental in numerical linear algebra [Parlett, 1998, Saad, 2011, Golub and Van Loan, 2013] and nonlinear optimization [Yuan, 2014, Liu et al., 2021c], specifically in the context of solving large linear systems and eigenvalue problems. In this section, we will briefly introduce the Krylov subspace and the Lanczos process, and recap some important properties; readers are referred to Saad [2011], Golub and Van Loan [2013] for more details.
@@
and the sequence of vectors $\left\{b,Ab,A^{2}b,\ldots,A^{N-1}b\right\}$ forms the basis for it. The Krylov subspace is widely acknowledged for its favorable properties in various aspects, including approximating eigenvalues [Kuczyński and Woźniakowski, 1992], solving the regularized nonconvex quadratic problems [Gould et al., 1999, Zhang et al., 2017, Carmon and Duchi, 2018], and reducing computation cost [Brown and Saad, 1990, Bellavia and Morini, 2001, Liu et al., 2013, Jiang et al., 2024].
The Lanczos process [Lanczos, 1950] is an algorithm that exploits the structure of the Krylov subspace when $A$ is symmetric. Specifically, in the $j$-th step of the Lanczos process, we can efficiently maintain an orthogonal basis $Q_{j}$ of $\mathcal{K}_{j}(A,b)$, so that $T_{j}=Q_{j}^{\top}AQ_{j}$ is tridiagonal, which means a tridiagonal matrix $T_{j}$ approximates $A$ in the Krylov subspace. Consequently, it allows to solve the minimal residual problem or the eigenvalue problem efficiently within the Krylov subspace. There are several equivalent variants of the Lanczos process [Paige, 1971, 1976, Meurant and Strakoš, 2006], and we follow the update rule as shown in Algorithm 3.
@@
We now present several key properties of the Krylov subspace and the Lanczos Process from Saad [2011].
@@
$A^{-1}v=-\frac{1}{c_{0}}\left(A^{m-1}v+c_{m-1}A^{m-2}v+\cdots+c_{2}Av+c_{1}v\right).$
@@
Denote the $n\times j$ matrix with column vectors $q_{1},\ldots,q_{j}$ by $Q_{j}$ and the $j\times j$ tridiagonal matrix by $T_{j}$, all of which are generated by Algorithm 3. Then the following three-term recurrence holds.
@@
Based on Proposition B.4, the Lanczos process is illustrated in Figure 7.
Figure 7: Classic three-term recurrence
@@ Appendix C Dynamic Lanczos Subroutine
This section lists the DLanczos subroutine (Algorithm 4) invoked in the LancBiO framework (Algorithm 2). One of the main differences between Algorithm 3 and Algorithm 4 is that Algorithm 3 represents the entire $m$-step Lanczos process, while Algorithm 4 serves as a one-step subroutine. Specifically, LancBiO invokes DLanczos once in each outer iteration (line 11 in Algorithm 2), expanding both $T$ and $Q$ by one dimension. Consequently, the inputs of Algorithm 4 are not indexed to avoid confusion, with their corresponding variables $(T_{k-1},Q_{k-1},A_{k},\beta_{k})$ in Algorithm 2 evolving across outer iterations indexed by $k$. Another difference lies in the dynamic property of the DLanczos subroutine, i.e., the matrix $A_{k}$ passed during each invocation varies in Algorithm 2, while the classic Lanczos process (Algorithm 3) employs a static $A$.
@@ Appendix D Extending LancBiO to Non-convex Lower-level Problem
The Lanczos process is known for its efficiency of constructing Krylov subspaces and is capable of solving indefinite linear systems [Greenbaum et al., 1999]. In this section, we will briefly demonstrate that the dynamic Lanczos-aided bilevel optimization framework, LancBiO, can also handle lower-level problems with the indefinite Hessian.
@@
with initial ponit $x_{0}$, initial residual $r_{0}=b-Ax_{0}$ and initial error $e_{0}=A^{-1}b-x_{0}$. If the matrix $A$ is positive-definite, the classic Lanczos algorithm is equivalent to the Conjugate Gradient (CG) algorithm [Hestenes et al., 1952], both of which minimize the $A$-norm of the error in an affine space [Greenbaum, 1997, Meurant and Strakoš, 2006], i.e., at the $m$-th step,
@@
If the matrix $A$ is not positive-definite, MINRES [Paige and Saunders, 1975] is the algorithm recognized to minimize the 2-norm of the residual in an affine space [Greenbaum, 1997, Meurant and Strakoš, 2006], i.e., at the $m$-th step,
@@
Additionally, based on $Q_{m}$ as the basis of the Krylov subspace $\mathcal{K}_{m}(A,b)$, $T_{m}$ as the projection of $A$ onto $\mathcal{K}_{m}(A,b)$, and the three-term recurrence
@@ Appendix E Proof of Smoothness of $y^{*}$ and $\varphi$
To ensure completeness, we provide detailed proofs for the preliminary lemmas that characterize the smoothness of the lower level solution $y^{*}$ and the hyper-objective $\varphi$.
@@
where $L_{\varphi}=\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left(L_{fx}+\frac{L_{gx}L_{fy}+L_{gxy}C_{fy}}{\mu_{g}}+\frac{L_{gx}C_{fy}L_{gyy}}{\mu_{g}^{2}}\right)$.
@@
$\left\|\left(A_{1}^{*}\right)^{-1}b_{1}-\left(A_{2}^{*}\right)^{-1}b_{2}\right\|\leq\frac{L_{fy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|+\frac{C_{fy}L_{gyy}}{\mu_{g}^{2}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|.$
In a similar way, the subsequent decomposition holds,
@@
$\left\|\nabla\varphi\left(x_{1}\right)-\nabla\varphi\left(x_{2}\right)\right\|\leq$ $~{}L_{fx}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|$
$+L_{gx}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left(\frac{L_{fy}}{\mu_{g}}+\frac{C_{fy}L_{gyy}}{\mu_{g}^{2}}\right)\left\|x_{1}-x_{2}\right\|$
$+L_{gxy}\frac{C_{fy}}{\mu_{g}}\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left\|x_{1}-x_{2}\right\|$
$=$ $~{}L_{\varphi}\left\|x_{1}-x_{2}\right\|,$
where $L_{\varphi}:=\left(1+\frac{L_{gx}}{\mu_{g}}\right)\left(L_{fx}+\frac{L_{gx}L_{fy}+L_{gxy}C_{fy}}{\mu_{g}}+\frac{L_{gx}C_{fy}L_{gyy}}{\mu_{g}^{2}}\right)$. ∎
@@
We rewrite the dynamic update rule from Section 2.2
@@
Note that $Q_{1}=\mathrm{span}\{q_{1}\}$ with $q_{1}=\frac{\bar{b}}{\left\|{\bar{b}}\right\|}$ satisfies (16). We will give a proof by induction. Suppose for $i=1,2,\ldots,j$, it holds that
@@
Although we can estimate the difference between the basis of the above two subspaces, it is noted that the Krylov subspaces can be very sensitive to small perturbation [Meurant and Strakoš, 2006, Paige, 1976, 1980, Greenbaum, 1997]. The next lemma interprets the perturbation analysis for the dynamic Lanczos process in terms of $A^{*}_{j}$, which satisfies an approximate three-term recurrence with a perturbation term $\delta Q$.
@@
Then, we consider the orthogonality of matrix $Q_{j}$, which is reflected by $R_{j}$ in (19). Multiply on both sides of (18) by $Q^{\top}_{j}$,
@@
Combining its symmetry with the decomposition (19), we obtain
@@
Denote $M_{j}=T_{j}R_{j}-R_{j}T_{j}$ which is upper triangular. Since the consecutive $q_{i}$ is orthogonal as revealed by (21), we conclude that the diagonal elements of $M_{j}$ are 00. Furthermore, by extracting the upper triangular part of the right hand side of (23), we can get
@@
From the boundedness of $\left\|{\delta q_{j}}\right\|$, it follows that for $1\leq s<t\leq j$, $\left|{\zeta_{st}}\right|\leq 2L_{gyy}\varepsilon_{j}$. ∎
Lemma F.2 illustrates the influence of the dynamics in Algorithm 2 imposed on the three-term Lanczos recurrence, and as (19) reveals, $R$ reflects the loss of orthogonality of the basis $Q$. However, the following lemmas demonstrate that the range of eigenvalues of the approximate projection matrix $T$ is indeed controllable.
To proceed, we establish the Ritz pairs of $T_{j}$ as $\left(\mu_{i}^{\left(j\right)},y_{i}^{\left(j\right)}\right)$ for $i=1,2,\dots,j$, such that
@@
The subsequent lemma describes the difference of eigenvalues between $T_{j}$ and some $T_{n}$ constructed in preceding steps.
@@
Multiply the extended eigenvector $y_{r}^{\left(i\right)}$ of $T_{i}$ by $T_{j}$, where $i<j$,
$T_{j}\left[\begin{array}[]{c}y_{r}^{\left(i\right)}\\ \mathbf{0}\\ \end{array}\right]=\left[\begin{array}[]{c}\mu_{r}^{\left(i\right)}y_{r}^{\left(i\right)}\\ \beta_{i+1}\varsigma_{ir}^{\left(i\right)}e_{1}\\ \end{array}\right].$ (25)
@@
and multiply the eigenvectors $\left(y_{s}^{\left(j\right)}\right)^{\top}$ and $y_{t}^{\left(j\right)}$ on the left and right of equation (20), respectively,
@@
Note that $\left|\epsilon_{st}^{\left(j\right)}\right|\leq 2jL_{gyy}\varepsilon_{j}$ because of Lemma F.2. Specifically, taking $s=t$ in (27),
@@
we can rewrite (29) in the matrix form
@@
it follows from the Cauchy–Schwarz inequality and (35) that
$\min_{1\leq d\leq l<j}\left|{\mu_{i}^{\left(j\right)}-\mu_{d}^{\left(l\right)}}\right|\leq\frac{\sqrt{j}\sqrt{\sum_{r=1}^{j-1}{\sum_{t=1}^{r}{\left(\epsilon_{tt}^{\left(r\right)}\right)^{2}}}}}{\left|\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}\,\leq\frac{2j^{2}L_{gyy}\varepsilon_{j}}{\sqrt{3}\left|\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}.$
@@
Suppose Assumptions 3.1 to 3.3 hold. $Q_{j}$ and $T_{j}$ are the basis matrix and the approximate tridiagonal matrix in the $j$-th step and $R_{j}$ is the strictly upper triangular matrix defined in (19) characterizing the orthogonality of $Q_{j}$. Given $\mu_{i}^{\left(j\right)}$ the $i$-th eigenvalue of $T_{j}$, then for $i=1,\ldots,j$,
@@
By conducting the left multiplication with $\left(y_{i}^{\left(j\right)}Q_{j}\right)^{\top}$ and the right multiplication with $y_{i}^{\left(j\right)}$ on both sides of equation (18), we obtain the following equation,
@@
Dividing it by $\left(z_{i}^{\left(j\right)}\right)^{\top}z_{i}^{\left(j\right)}$, we have
$\left|\frac{\left(z_{i}^{\left(j\right)}\right)^{\top}A_{j}^{*}z_{i}^{\left(j\right)}-\mu_{i}^{\left(j\right)}\left(z_{i}^{\left(j\right)}\right)^{\top}z_{i}^{\left(j\right)}}{\left(z_{i}^{\left(j\right)}\right)^{\top}z_{i}^{\left(j\right)}}\right|$ $\leq\frac{\left|\epsilon_{ii}^{\left(j\right)}\right|+\left|\left(z_{i}^{\left(j\right)}\right)^{\top}\delta Q_{j}y_{i}^{\left(j\right)}\right|}{\left|1+2\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|}$
$\leq\frac{\left|\epsilon_{ii}^{\left(j\right)}\right|+L_{gyy}\varepsilon_{j}\sqrt{j\left(\left|1+2\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|\right)}}{\left|1+2\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}\right|},$ (37)
where the inequalities come from
@@
$\left|{\left(y_{i}^{\left(j\right)}\right)^{\top}R_{j}y_{i}^{\left(j\right)}}\right|<\frac{3}{8}-\frac{\varepsilon_{j}}{2}$ (39)